# Analytical results for the distribution of first return times of non-backtracking random walks on configuration model networks

### v1

@@ 5 Summary
We presented analytical results for the distribution of first return times of NBWs on configuration model networks consisting of $N$ nodes with degree distribution $P(k)$, focusing on the case in which the network consists of a single connected component. It was found that the tail distribution $P(T_{\rm FR}>t)$ of first return times is given by a discrete Laplace transform of the degree distribution $P(k)$. This result demonstrates the relation between structural properties of a network, captured by the degree distribution, and the properties of dynamical processes taking place on the network. It was found that $P(T_{\rm FR}>t)$ exhibits an exponential tail, which is determined by the properties of the low-degree nodes that reside in the periphery of the network. We calculated the mean first return time and found that ${\mathbb{E}}[T_{\rm FR}]=\langle\frac{Nc}{K}\rangle$. Surprisingly, this result coincides with the result of the Kac’s lemma that applies to classical RWs, in agreement with recent rigorous results by Fasino et al. [27]. We also calculated the variance ${\rm Var}(T_{\rm FR})$, which accounts for the variability of the first return times between different NBW trajectories. We applied this formalism to random regular graphs, Erdős-Rényi networks and configuration model networks with exponential and power-law degree distributions and obtained closed-form expressions for $P(T_{\rm FR}>t)$ and its first two moments. These results provide useful insight on the advantages of NBWs over classical RWs in network exploration, sampling and search processes. Our results are expected to hold for a broader class of networks, in which the mixing time is much shorter than the mean first return times.
This work was supported by Grant no. 2020720 from the United States –Israel Binational Science Foundation (BSF) and grant no. 2102832 from the National Science Foundation (NSF).

### v2

@@ 7 Summary
We presented analytical results for the distribution of first return times of NBWs on configuration model networks consisting of $N$ nodes with degree distribution $P(k)$, focusing on the case in which the network consists of a single connected component. It was found that the tail distribution $P(T_{\rm FR}>t)$ of first return times is given by a discrete Laplace transform of the degree distribution $P(k)$. This result demonstrates the relation between structural properties of a network, captured by the degree distribution, and the properties of dynamical processes taking place on the network. It was found that $P(T_{\rm FR}>t)$ exhibits an exponential tail, which is determined by the properties of the low-degree nodes that reside in the periphery of the network. We calculated the mean first return time and found that ${\mathbb{E}}[T_{\rm FR}]=\langle\frac{Nc}{K}\rangle$. Surprisingly, this result coincides with the result of Kac’s lemma that applies to simple RWs, in agreement with recent rigorous results by Fasino et al. [30]. We also calculated the variance ${\rm Var}(T_{\rm FR})$, which accounts for the variability of the first return times between different NBW trajectories. We applied this formalism to random regular graphs, Erdős-Rényi networks and configuration model networks with exponential and power-law degree distributions and obtained closed-form expressions for $P(T_{\rm FR}>t)$ and its first two moments. These results provide useful insight on the advantages of NBWs over simple RWs in network exploration, sampling and search processes. Our results are expected to hold for a broader class of networks, in which the mixing time is much shorter than the mean first return times.
This work was supported by Grant no. 2020720 from the United States-Israel Binational Science Foundation (BSF) and grant no. 2102832 from the National Science Foundation (NSF).